# Disentangling the anisotropic radio sky: Fisher forecasts for 21cm arrays

### v1

@@ Abstract
The existence of a radio synchrotron background (RSB) is implied by a number of measurements, including excess emission seen by the ARCADE2 and LWA experiments. Highly sensitive wideband radio arrays, of the kind used to measure the cosmic 21-cm signal, provide a promising way to further constrain the RSB through its anisotropy, providing additional insight into its origin. We present a framework for evaluating the potential of 21-cm arrays to disentangle different components of the diffuse radio sky based on the combination of their frequency spectrum and angular power spectrum (APS). The formalism is designed to calculate uncertainties due to the intrinsic cosmic variance alone or together with instrumental noise. In particular, we predict the potential for measuring the anisotropy of a broad generalised class of excess radio background models using the low-frequency HERA array as an example. We find that a HERA-like array can distinguish an RSB from other sky components based on its angular clustering and spectral dependence, even if these are quite similar to one or more of the other components – but only in the case that the RSB is relatively bright.
@@ 1 Introduction
The bright diffuse radio emission at high Galactic latitudes has long been known and has been observed in a number of observations (e.g. Westerhout & Oort, 1951). For a long time, in the absence of good constraints on both the Galactic emission and the line-integrated emission from extragalactic point sources, typical analyses simply assumed that the diffuse radio emission was mainly a combination of Galactic emission and extragalactic point sources. This understanding has led to a lack of motivation to consider any significant radio background, which may be a significant oversight of 21-cm cosmology, i.e. the search for the neutral hydrogen lines embedded in the primordial light from Recombination. Recent re-examinations of the composition of the diffuse radio sky were triggered by ARCADE 2 (Fixsen et al., 2011), which used two empirical galactic foreground models to show that, in addition to the primordial $\sim$3 K blackbody radiation, there is an excess power-law spectrum in the monopole of the diffuse radio background (Dowell & Taylor, 2018),
@@
This excess temperature is also known as the Radio Synchrotron Background (RSB).
@@
The origin of the RSB has become one of the major puzzles in contemporary astrophysics, since both the known galactic processes and the known classes of extragalactic point sources have difficulty in explaining such a large excess (see Singal et al. (2023) for a recent review). A number of exotic models have been proposed. Some works have considered the modified population models of faint, unresolved point sources (e.g. Vernstrom et al., 2014; Condon et al., 2012; Hardcastle et al., 2021). Biermann et al. (2014) investigated the possibility that supernovae of massive population \@slowromancapiii@ stars are the source of the diffuse background. Some other exotic astrophysical processes have also been considered, such as injections from high energy particles (Cline & Vincent, 2013), emission from stellar black holes (Ewall-Wice et al., 2018), primordial black holes (Mittal & Kulkarni, 2022), and others more related to structures on cosmological scales, such as annihilating dark matter (DM) in halos or filaments (Fortes et al., 2019), superconducting cosmic strings (Cyr et al., 2023).
Strong observational tests are needed for these candidate theoretical models. Theoretically, different models have different predictions for RSB random field statistics. The rich information can be extracted from its frequency-frequency angular power spectrum, which in turn can help to filter out these exotic models. However, appropriate analysis frameworks are currently lacking. Although the analyses of Offringa et al. (2022) and Cowie et al. (2023) were designed to study the anisotropy of the excess background, they simply assume that the RSB dominates the fluctuations in the high-latitude radio sky, and therefore do not separate out the galactic components. To some extent, dedicated statistical measurements of the frequency and spatial structure of the RSB field are still needed, especially for an RSB identification strategy that is logically independent of the ARCADE 2 result and its foreground subtraction strategies.
@@
In this work, to encapsulate our ignorance of each component, we model the diffuse sky as the sum of several independent random fields and constrain them all together. We use a generalised parametric form to describe the angular and frequency structures of the random fields. In this way, each field is specified by an independent set of parameters. The statistical model is described in terms of spherical harmonic (SH) modes whose statistics are characterised by a multivariate Gaussian distribution specified by a parameterised angular power spectrum and frequency-frequency covariance matrix. Assuming a likelihood-based joint analysis for all fields, we provide a formalism based on Fisher matrix techniques for the prediction of constraints on the RSB model using low-frequency radio interferometric arrays. The function and purpose of this formalism is twofold: (1) to predict whether the intrinsic uncertainty caused by the cosmic variance allows us to disentangle different components, given that we do not have an ensemble of universes to use, but finite SH modes of the field configurations; (2) to predict whether the experimental setup allows us to detect RSBs and discriminate between different predictions. These are the two things we need to do in order to design an experiment to test particular RSB models.
@@
Using this framework, we present a Fisher forecast for the Hydrogen Epoch of Reionization Array (HERA; DeBoer et al., 2017) to measure the RSB angular power spectrum (APS). Without going into specific theoretical models, we assume three different APS models and analyse the tightness of HERA’s constraints on the RSB parameters for each of the cases, without imposing any prior on any parameter of any component of the radio sky.
@@
The rest of the paper is structured as follows: In Section 2, we discuss the basic formalism of Fisher analysis, including a general statistical model (Sec. 2.1) and the intrinsic uncertainty for constraining the statistical models of the random fields (Sec. 2.2). In Section 2.3, we present a formalism for predicting the ability of the 21 cm radio interferometric array to separate different components of the radio sky. In Section 3 we give a specific example where we analyse the ability of the HERA array to detect RSBs. In Section 4 we discuss and conclude this work.
@@ 2 Fisher Forecast Formalism
Our formalism is based on Fisher matrix techniques and assumes that the diffuse radio sky can be represented as a few statistically independent components, for each of which the distribution of SH modes is approximated by a multivariate complex Gaussian distribution. This formalism is designed to predict the accuracy of 21-cm arrays for disentangling the anisotropic radio sky. In particular, it can be used to predict their ability to test exotic models for the RSB.
@@ 2.1 Statistical model
where the superscript ‘s’ indexes the different sky components. Typically, our understanding of the sky is in terms of different types of emission sources, such as galactic synchrotron emission, galactic free-free emission, the RSB, etc.
@@
where $A$ denotes the magnitude of the power spectrum at the reference frequency $\nu_{\rm ref}$ and the reference scale $\ell_{\rm ref}$. $\alpha$ describes the spatial structure of the random field: the power spectrum of a naturally occurring random field is usually smooth enough that we can approximate it as a power law over a modest interval of $\ell$. $\beta$ and $\xi$ describe the frequency structure of the radio emission, where $\beta$ describes a power-law approximation of the spectrum and $\xi$ describes the degree of anticorrelation, which can be thought of as a low-order correction to the power law. We have a set of parameters $\{A^{(s)},\alpha^{(s)},\beta^{(s)},\xi^{(s)}\}$ for each component field indexed by $s$.
2.2 Cosmic variance and component discriminability
We approximate the distribution of spherical harmonic modes as a multivariate complex normal ($\mathcal{CN}$) distribution. For radio interferometric arrays, if we only consider crosscorrelation visibilities, the measurement is basically only sensitive to fluctuating modes and has no response to the monopole. We will therefore only consider the zero mean, $\ell>0$ modes,
@@ 2.2 Cosmic variance and component discriminability
Since we do not really have an ensemble of universes, but finite spherical harmonics of the field configurations in our universe. Therefore, before testing the angular power spectrum predicted by the model, we need to know whether the cosmic variance allows us to separate the RSB from other components. In other words, the challenge posed by the strategy of constraining all fields simultaneously is whether we can tell the statistical differences between different fields using a finite set of spherical harmonics; if it does not, we are forced to use priors or adapt the experimental setup to look at more frequencies or more spherical harmonic modes.
@@
In order to better discuss the effect of the cosmic variance on the model constraints, we perform a simple set of comparative analyses. We assume a radio sky with four statistically independent components; Galactic synchrotron, Galactic free-free, Extragalactic point sources and Extragalactic free-free. Their fiducial parameter values are given in Section 3. We compare how tightly different SH combinations, as listed in Table 1, constrain these parameters using equation (10). The results of the Fisher forecasts, the simulated measurements and the uncertainties are shown in Table 2, and the uncertainties of each parameter are visualised in Fig. 1. Comparisons between Forecasts Low res., 21 freqs. and Wideband show that, in general, more spherical harmonic modes at more frequencies provide tighter constraints. However, comparing Forecasts 21 freqs., 101 freqs. and Wideband also shows that it is not always the case that a higher resolution of frequencies is better. This can be interpreted as a consequence of the fact that all field components in the model are more or less highly correlated in frequency. If the difference between two frequencies is much smaller than the correlation length, then the pair will provide no more information than a single frequency. Even in extreme cases, too high a frequency resolution can cause the frequency-frequency covariance matrix to become numerically singular and therefore not accurately inverted.
@@ 2.3 Fisher forecast for radio interferometric arrays
The evaluation of the intrinsic uncertainty due to the cosmic variance is a prerequisite for component separation. Since the purpose of component separation may vary, the impact of the intrinsic uncertainty caused by the cosmic variance should be evaluated in an ad hoc manner. In this analysis we assume that the cosmic variance is not a limit for constraining the RSB and other dominant components of the diffuse sky with a particular set of spherical harmonics and frequencies. The next question is whether a 21 cm array, in particular its noise, spherical harmonic response and observing strategy, will allow us to separate these components accurately.
@@
where we have applied the Woodbury matrix identity. After applying some algebra tricks, equation (16) can be rewritten as
@@ 3.1 Experimental Setup
The HERA 21-cm array is composed of 14 m parabolic dishes with wide-band Vivaldi feeds suspended at prime focus. We simulate HERA Band 1 (roughly from 115 MHz to 135 MHz) observations in a relatively radio quiet part of the survey region: Field C (RA: 4.0–6.25 hours).
@@ 3.2 Fiducial models
RSB 1 $(6.63\text{ K})^{2}$ $-2.4$ $-2.66$ $4.0$
RSB 2 $(2.10\text{ K})^{2}$ $-3.0$ $-2.66$ $4.0$
RSB 3 $(21\text{ K})^{2}$ $0.0$ $-2.66$ $4.0$
@@
Figure 3: The assumed angular power spectra of different components of the radio sky at the reference frequency of 130 MHz. Note that the RSB1 model overlaps the Galactic synchrotron component. The opacity of the gray background is proportional to the overall response of the HERA array model used in this work to $C_{\ell}$. The overall $C_{\ell}$ response, which is defined as $\sum_{t,m,\boldsymbol{b}_{ij}}|X(\boldsymbol{b}_{\rm ij},\nu_{\rm ref},t,\ell,m)|^{2}/\sigma_{\rm ij}^{2}(t)$, characterizes the linear mapping from $C_{\ell}$ to the whitened data covariance, $\langle\boldsymbol{d}^{\dagger}\mathbf{N}^{-1}\boldsymbol{d}\rangle$. For computational reasons we have only calculated up to $\ell_{\rm max}=90$.
We assume that the radio sky consists mainly of five statistically independent components, namely galactic synchrotron, galactic free-free, extragalactic point sources, extragalactic free-free and the radio synchrotron background. For the first four, we use the parameters in Santos et al. (2005) as the fiducial values for the input to the Fisher analysis (see Table 3 and Fig. 3).
@@
For the RSB, since there are currently no well-motivated fiducial values other than the frequency power law index, i.e. $\beta\simeq-2.66$. We consider three simple phenomenological models, using the Galactic Synchrotron as a reference, to test the ability of HERA to constrain the RSB model parameters under different circumstances:
  • RSB 1: The parameters, except for $\beta$, are the same as those of the Galactic synchrotron. This model mimics the galactic synchrotron, but differs from it only in its spectral structure.
  • RSB 2: At the reference mode ($\ell_{\rm ref}=10,\nu_{\rm ref}=130\text{ MHz}$), the RSB is an order of magnitude stronger than the Galactic synchrotron, but has a flatter angular power law.
@@
  • RSB 3: At the reference mode ($\ell_{\rm ref}=10,\nu_{\rm ref}=130\text{ MHz}$), the RSB is an order of magnitude less than the Galactic synchrotron, but has a steeper angular power law.
@@
Theoretically, we could have also considered a higher (weaker) power and a steeper (flatter) angular power law. However, such an assumption implies that the RSB monopole is much stronger (weaker) than the Galactic synchrotron monopole, considering a brutal extrapolation of the APS model to the $\ell=0$ end. Such low $\ell$ structures do not seem to be supported by the current monopole measurements. Also, as an illustrative example, we do not take the trouble to exhaust all the other possible models.

### v2

@@ Abstract
The existence of a radio synchrotron background (RSB) excess is implied by a number of measurements, including excess emission seen by the ARCADE 2 and LWA experiments. Highly sensitive wideband radio arrays, of the kind used to measure the cosmic 21cm signal, provide a promising way to further constrain the RSB excess through its anisotropy, providing additional insight into its origin. We present a framework for evaluating the potential of 21cm arrays to disentangle different components of the diffuse radio sky based on the combination of their frequency spectrum and angular power spectrum (APS). The formalism is designed to calculate uncertainties due to the intrinsic cosmic variance alone or together with instrumental noise. In particular, we predict the potential for measuring the anisotropy of a broad generalised class of excess radio background models using the low-frequency HERA array as an example. We find that a HERA-like array can distinguish an RSB excess from other sky components based on its angular clustering and spectral dependence, even if these are quite similar to one or more of the other components – but only in the case that the RSB excess is relatively bright.
@@ 1 Introduction
The bright diffuse radio emission at high Galactic latitudes has long been known and has been observed in a number of observations (e.g. Westerhout & Oort, 1951). For a long time, in the absence of good constraints on both the Galactic emission and the line-integrated emission from extragalactic point sources, typical analyses simply assumed that the diffuse radio emission was mainly a combination of Galactic emission and extragalactic point sources. This understanding has led to a lack of motivation to consider any significant radio background, which may be a significant oversight of 21cm cosmology, i.e. the search for the neutral hydrogen lines embedded in the primordial light from Recombination. Recent re-examinations of the composition of the diffuse radio sky were triggered by ARCADE 2 (Fixsen et al., 2011), which used two empirical Galactic foreground models to show that, in addition to the primordial $\sim$3 K blackbody radiation, there is an excess power-law spectrum in the monopole of the diffuse radio background (Dowell & Taylor, 2018),
@@
This excess temperature is also known as the Radio Synchrotron Background (RSB). Although Synchrotron in this context refers to the power law determined by ARCADE 2, it does not necessarily imply an origin exclusively from synchrotron emission. Moreover, given the unknown redshifts of the origins of the RSB excess, the terms “foreground” and “background” are no longer unambiguously defined in 21cm cosmology. To avoid the confusion that these terms may cause, we hereby explicitly divide the radio sky into five segments: Galactic synchrotron, Galactic free-free, Extragalactic free-free, Extragalactic point sources, and RSB excess (excluding, for simplicity, the subdominant 21cm emission). Note that since extragalactic point sources models or catalogues are usually defined up to a low flux limit, the definition of the RSB excess should not only include unknown exotic contributions (if they exist), but also potential residual extragalactic point sources.
@@
The origin of the RSB excess has become one of the major puzzles in contemporary astrophysics, since both the known Galactic processes and the known classes of extragalactic point sources have difficulty in explaining such a large excess (see Singal et al. (2023) for a recent review). A number of exotic models have been proposed. Some works have considered the modified population models of faint, unresolved point sources (e.g. Vernstrom et al., 2014; Condon et al., 2012; Hardcastle et al., 2021). Biermann et al. (2014) investigated the possibility that supernovae of massive population \@slowromancapiii@ stars are the source of the diffuse background. Some other exotic astrophysical processes have also been considered, such as injections from high energy particles (Cline & Vincent, 2013), emission from stellar black holes (Ewall-Wice et al., 2018), primordial black holes (Mittal & Kulkarni, 2022), and others more related to structures on cosmological scales, such as annihilating dark matter (DM) in halos or filaments (Fortes et al., 2019), superconducting cosmic strings (Cyr et al., 2023).
Strong observational tests are needed for these candidate theoretical models. Theoretically, different models have different predictions for the random field statistics of the RSB excess. A rich source of information is the frequency-frequency angular power spectrum, which in turn can help to filter out these exotic models. Some recent work has begun to recognise this, e.g. Todarello et al. (2024) suggests a 20% lower bound on the extragalactic contribution to the RSB (which in this paper should be understood as the sum of the RSB excess, the extragalactic point sources, and the extragalactic free-free emission) by studying the angular cross-correlation of LOFAR images of the diffuse radio sky with matter tracers at different redshifts provided by galaxy catalogues and CMB lensing. Offringa et al. (2022) and Cowie et al. (2023) investigated the anisotropy of the RSB by examining high multipole modes within fields at high Galactic latitudes. The former analysis attempted to identify a significant Galactic component within the probed range of $\ell$, but concluded that such a component was absent; the latter suggests that while the anisotropy power for $\ell<7000$ is predominantly due to diffuse Galactic emission, this is not the case for $\ell>7000$, as indicated by the change in the power law at $\ell\approx 7000$. Nevertheless, there remains a gap in current research methodologies to effectively isolate the RSB excess. Dedicated statistical measurements of the frequency and spatial structure of the RSB excess are essential to refine our understanding of the origins of the RSB excess, especially for an RSB excess separation strategy that is logically independent of the ARCADE 2 result and its foreground subtraction strategies.
@@
In this work, to encapsulate our ignorance of each component, we model the diffuse sky as the sum of several independent random fields and constrain them all together. We use a generalised parametric form to describe the angular and frequency structures of the random fields. In this way, each field is specified by an independent set of parameters. The statistical model is described in terms of spherical harmonic (SH) modes whose statistics are characterised by a multivariate Gaussian distribution specified by a parameterised angular power spectrum and frequency-frequency covariance matrix. Assuming a likelihood-based joint analysis for all fields, we provide a formalism based on Fisher matrix techniques for the prediction of constraints on the RSB excess model using low-frequency radio interferometric arrays. The function and purpose of this formalism is twofold: (1) to predict whether the intrinsic uncertainty caused by the cosmic variance allows us to disentangle different components, given that we do not have an ensemble of universes to use, but finite SH modes of the field configurations; (2) to predict whether the experimental setup allows us to detect the RSB excess and discriminate between different predictions. These are the two things we need to do in order to design an experiment to test particular RSB excess models.
@@
Using this framework, we present a Fisher forecast for the Hydrogen Epoch of Reionization Array (HERA; DeBoer et al., 2017) to measure the angular power spectrum (APS) of the RSB excess. Without going into specific theoretical models, we assume three different APS models and analyse the tightness of HERA’s constraints on the RSB excess parameters for each of the cases, without imposing any prior on any parameter of any component of the radio sky.
@@
The rest of the paper is structured as follows: In Section 2, we discuss the basic formalism of Fisher analysis, including a general statistical model (Sec. 2.1) and the intrinsic uncertainty for constraining the statistical models of the random fields (Sec. 2.2). In Section 2.3, we present a formalism for predicting the ability of the 21cm radio interferometric array to separate different components of the radio sky. In Section 3 we give a specific example where we analyse the ability of the HERA array to detect the RSB excess. In Section 4 we discuss and conclude this work.
@@ 2 Fisher Forecast Formalism
Our formalism is based on Fisher matrix techniques and assumes that the diffuse radio sky can be represented as a few statistically independent components, for each of which the distribution of SH modes is approximated by a multivariate complex Gaussian distribution. This formalism is designed to predict the accuracy of 21cm arrays for disentangling the anisotropic radio sky. In particular, it can be used to predict their ability to test exotic models for the RSB excess.
@@ 2.1 Statistical model
where the superscript ‘$s$’ indexes the different sky components. Typically, our understanding of the sky is in terms of different types of emission sources, such as Galactic synchrotron emission, Galactic free-free emission, one or several RSB excess emissions, etc.
@@
where $A$ denotes the magnitude of the power spectrum at the reference frequency $\nu_{\rm ref}$ and the reference scale $\ell_{\rm ref}$. $\alpha$ describes the spatial structure of the random field: the power spectrum of a naturally occurring random field is usually smooth enough that we can approximate it as a power law over a modest interval of $\ell$. $\beta$ and $\xi$ describe the frequency structure of the radio emission, where $\beta$ describes a power-law approximation of the spectrum and $\xi$ describes the degree of anticorrelation, which can be thought of as a low-order correction to the power law. We have a set of parameters $\{A^{(s)},\alpha^{(s)},\beta^{(s)},\xi^{(s)}\}$ for each component field indexed by $s$. The goal of this analysis is to determine whether the values of the parameters $A$, $\alpha$, $\beta$, and $\xi$ for the RSB excess could be constrained by a suitable 21cm interferometer experiment.
2.2 Cosmic variance and component distinguishability
We approximate the distribution of spherical harmonic modes as a multivariate complex normal ($\mathcal{CN}$) distribution. For radio interferometric arrays, if we only consider cross-correlation visibilities, the measurement is basically only sensitive to fluctuating modes and has no response to the monopole. We will therefore only consider the zero mean, $\ell>0$ modes,
@@ 2.2 Cosmic variance and component distinguishability
We do not really have an ensemble of universes, but finite spherical harmonics of the field configurations in our Universe. Therefore, before testing the angular power spectrum predicted by the model, we need to know whether the cosmic variance allows us to separate the RSB excess from other components. In other words, the challenge posed by the strategy of constraining all fields simultaneously is whether we can tell the statistical differences between different fields using a finite set of spherical harmonics; if it does not, we are forced to use priors or adapt the experimental setup to look at more frequencies or more spherical harmonic modes.
@@
In order to better discuss the effect of the cosmic variance on the model constraints, we first perform a simple set of comparative analyses. We assume a radio sky with four statistically-independent components; two Galactic components (Galactic synchrotron and Galactic free-free), and two RSB components (Extragalactic point sources and Extragalactic free-free). In presenting an illustrative example of our cosmic variance evaluation scheme, we choose not to include a potential exotic RSB excess. This represents the most optimistic case for a statistical model of the anisotropic sky, in which we have a complete model with a reasonably minimal number of statistical parameters. Adding an exotic component (as we shall do in the next section) necessarily increases the number of parameters, and the structure and even the amplitude of the excess is not well constrained. The fiducial parameter values are given in Section 3. We compare how tightly different SH combinations, as listed in Table 1, constrain these parameters using equation (10). The results of the Fisher forecasts, the simulated measurements and the uncertainties are shown in Table 2, and the uncertainties of each parameter are visualised in Fig. 1. Comparisons between Forecasts Low res., 21 freqs. and Wideband show that, in general, more spherical harmonic modes at more frequencies provide tighter constraints. However, comparing Forecasts 21 freqs., 101 freqs. and Wideband also shows that it is not always the case that a higher resolution of frequencies is better. This can be interpreted as a consequence of the fact that all field components in the model are more or less highly correlated in frequency. If the difference between two frequencies is much smaller than the correlation length, then the pair will provide no more information than a single frequency. Even in extreme cases, a frequency resolution that is too high can cause the frequency-frequency covariance matrix to become numerically singular and therefore not accurately inverted.
@@ 2.3 Fisher forecast for radio interferometric arrays
The evaluation of the intrinsic uncertainty due to the cosmic variance is a prerequisite for component separation. Since the purpose of component separation may vary, the impact of the intrinsic uncertainty caused by the cosmic variance should be evaluated in an ad hoc manner. In this analysis we assume that the cosmic variance is not a limit for constraining the RSB excess and other dominant components of the diffuse sky with a particular set of spherical harmonics and frequencies. The next question is whether a 21cm array, in particular its noise, spherical harmonic response and observing strategy, will allow us to separate these components accurately.
@@
where we have applied the Woodbury matrix identity. After some algebra steps, equation (16) can be rewritten as
@@ 3.1 Experimental Setup
The HERA 21cm array is composed of 14 m parabolic dishes with wide-band Vivaldi feeds suspended at prime focus. We simulate HERA Band 1 (roughly from 115 MHz to 135 MHz) observations in a relatively radio quiet part of the survey region: Field C (RA: 4.0–6.25 hours).
@@ 3.2 Fiducial models
RSB excess 1 $(6.63\text{ K})^{2}$ $-2.4$ $-2.66$ $4.0$
RSB excess 2 $(2.10\text{ K})^{2}$ $-3.0$ $-2.66$ $4.0$
RSB excess 3 $(21\text{ K})^{2}$ $0.0$ $-2.66$ $4.0$
@@
Figure 3: The assumed angular power spectra of different components of the radio sky at the reference frequency of 130 MHz. Note that the RSB excess 1 overlaps the Galactic synchrotron component. The opacity of the shaded background is proportional to the overall response of the HERA array model used in this work to $C_{\ell}$. The overall $C_{\ell}$ response, which is defined as $\sum_{t,m,\boldsymbol{b}_{ij}}|X(\boldsymbol{b}_{\rm ij},\nu_{\rm ref},t,\ell,m)|^{2}/\sigma_{\rm ij}^{2}(t)$, characterizes the linear mapping from $C_{\ell}$ to the whitened data covariance, $\langle\boldsymbol{d}^{\dagger}\mathbf{N}^{-1}\boldsymbol{d}\rangle$. For computational reasons we have only calculated up to $\ell_{\rm max}=90$.
We assume that the radio sky consists mainly of five statistically independent components, namely Galactic synchrotron, Galactic free-free, extragalactic point sources, extragalactic free-free and the radio synchrotron background. For the first four, we use Eq. 6 and the parameter values that were presented in Santos et al. (2005) as the fiducial values for the Fisher analysis (see Table 3 and Fig. 3), except for the Galactic synchrotron spectral index, for which we use $\beta=-2.6$. This is a typical value for high Galactic latitudes between 90 and 190 MHz (Padovani et al., 2021). This fiducial Galactic synchrotron power law is flatter than the fiducial RSB power law, reflecting the current perception in the RSB literature (e.g. Dowell & Taylor, 2018) that the RSB excess plays a more important role at lower frequencies. Note also the chosen spectral index of $-2.07$ for the extragalactic point sources used in Santos et al. (2005). This flattened spectrum was obtained by first extrapolating the spectral index in the low frequency interval down to 0.1 mJy, the low flux cutoff below which the clustering signal dominates over the Poisson signal. It was then extrapolated to the spectral index in the lower frequency interval based on the observation that the source spectra flattened from the ‘low’ frequency interval of 327 MHz to 1.4 GHz to the ‘very low’ frequency interval of 74 MHz to 327 MHz, with a median change in the spectral index of $\Delta\beta=0.24$.
@@
For the RSB excess, since there are currently no well-motivated fiducial values other than the frequency power law index, i.e. $\beta\simeq-2.66$, we consider three simple phenomenological models, using the Galactic Synchrotron as a reference, to test the ability of HERA to constrain the RSB excess parameters under different circumstances:
  • RSB excess 1: The parameters, except for $\beta$, are the same as for Galactic synchrotron. This model mimics the Galactic synchrotron in terms of its angular power spectrum and amplitude, differing from it only in its frequency structure. We include this as a toy model for an RSB excess that would be difficult to distinguish from foreground emission based on angular clustering statistics alone.
@@
  • RSB excess 2: At the reference point ($\ell_{\rm ref}=10,\nu_{\rm ref}=130\text{ MHz}$), the anisotropy power of the RSB is an order of magnitude less than the Galactic synchrotron, but has a steeper angular power law.
  • RSB excess 3: At the reference point ($\ell_{\rm ref}=10,\nu_{\rm ref}=130\text{ MHz}$), the anisotropy power of the RSB is an order of magnitude stronger than that of the Galactic synchrotron, but has a flatter angular power law.
@@
The RSB excess 1 is not physically motivated and we discuss the discriminability of the model more in a numerical sense, as proximity in parameter values could lead to degeneracies in the model, where different combinations of parameter values could lead to similar predictions. On the other hand, RSB excess 2 has a steeper APS where the larger (angular) scale fluctuations dominate. RSB excess 3 is an example of a flatter APS, where smaller (angular) scale fluctuations play a more important role. This particular model represents white noise extragalactic emission. All three of these cases are intended to be toy models that represent phenomenological scenarios, and are not intended to mimic particular physical explanations for the excess. Theoretically, we could have also considered a higher (weaker) power and a steeper (flatter) angular power law. However, such an assumption implies that the monopole of the RSB excess is much stronger (weaker) than the Galactic synchrotron monopole, considering a brutal extrapolation of the APS model to the $\ell\to 0$ end. Such low $\ell$ structures do not seem to be supported by the current monopole measurements. Also, as an illustrative example, we do not take the trouble to exhaust all the other possible models.